1. A duly-licensed engineer designs two houses that are shaped and positioned like a part of the branches of the hyperbola whose equation is 625𝑦2 − 400𝑥2 = 250,000, where x and y are in yards. How far apart are the houses at

their closest point?


1
Expert's answer
2021-10-25T16:40:18-0400
"625\ud835\udc66^2 \u2212 400\ud835\udc65^2 = 250000"

"\dfrac{y^2}{400}-\dfrac{x^2}{625}=1"

Vertices: "(0, -20), (0, 20)."




The distance between the closest points is


"20-(-20)=40(yards)"

The houses at their closest point are 40 yards apart.